OpenStudy (anonymous):

Find d/dx (x^6/30? Is the answer x^5/5

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Just remember that (x^5)/5 would be a less confusing markup online.

OpenStudy (anonymous):

Thanks!

OpenStudy (unklerhaukus):

\[{d\over dx} {x^6 \over 30}\]\[={1 \over 30}{d\over dx} x^6\]\[={1 \over 30}6x^5\]\[={x^5 \over 5}\]

OpenStudy (anonymous):

I envy your command of LaTeX.